The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu, Dominik J. Bomans, and Eva K. Grebel

                                 mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                 http://www.astro.uni-wuerzburg.de/~grebel/MCNews.html

No. 10 Apr 10, 1997

Table of Contents

News and Views  1
Abstracts of 6 refereed papers  2
Abstract of 1 non-refereed paper 6
Job Advertisements  7
Meeting Announcements  8

News and Views

Widening the Scope of MCNews?

In the past issue of the Magellanic Clouds Newsletter, we solicited opinion about widening the scope of the Newsletter to include also Magellanic irregulars. We received a number of comments ranging from enthusiastic endorsement to rejection of the idea. Taking all arguments into account we have decided to confine the Newsletter to the Magellanic Clouds for the time being.

Where are the Ring Nebulae around B Supergiants?

The rings around SN 1987A raise an intriguing question: how common are such rings around earlytype B supergiants? We have examined ∼60 early-type B supergiants in the Magellanic Clouds or in the Galaxy, and find no evidence for any ring nebulae. Our experience cannot have been unique; other people must have made similar searches that yielded negative results.

Negative results for a small sample of stars are not likely to be reported in the literature. Meanwhile, the lack of published reports may stimulate more searches, which would produce more negative results that will never be published. To prevent duplicating effort and telescope time, we propose that those who have obtained negative results in the search of rings around B supergiants pool their data together to assemble a statistically meaningful sample for further analysis. Useful data include long-slit spectra of early B supergiants in the Magellanic Clouds and Hα images of early B supergiants in the Galaxy. Those who are interested please contact one of us listed below. Thank you.

You-Hua Chu (chu@astro.uiuc.edu) Wolfgang Brandner (brandner@astro.uiuc.edu) Eva K. Grebel (grebel@astro.uni-wuerzburg.de)

Paper Abstracts

The Luminous Eclipsing SMC OB + WN Binary HD 5980 Before and During the Recent LBV-Like Outburst: An Extreme Case of Colliding Winds

A.F.J. Moffat^{1,2,3}, S.V. Marchenko^{1,4}, P. Bartzakos^{1,3}, V.S. Niemela^{5,6,3}, M.A. Cerruti^{7,8,3}, A.M. Magalhaes^{8}, L. Balona^{9}, N. St-Louis^{1}, W. Seggewiss^{10,2}, and R. Lamontagne^{1}

1 Département de Physique, Université de Montréal, C.P. 6128, Succ. Centre-Ville, Montréal, QC, H3C 3J7, Canada, and Observatoire du mont Mégantic 2 Visiting Astronomer, Max Planck Institut für Astronomie 2.2m telescope, La Silla, Chile 3 Visiting Astronomer, CASLEO, operated under agreement between CONICET, SeCyt, and the Universities of La Plata, Córdoba and San Juan, Argentina 4 Visiting Astronomer, Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatories, operated by the Association of Universities for Research in Astronomy, Inc., under cooperative agreement with the National Science Foundation 5 Facultad de Ciencias Astronónicas y Geofísicas, Universidad Nacional de La Plata, 1900 La Plata, Argentina 6 Member of the Carrera del Investigador Científico, CIC, Prov. de Buenos Aires, Argentina 7 Member of the Carrera del Investigador Científico, CONICET, Argentina 8 Inst. Astron. e Geofisico, Univ. de Sao Paulo, Caixa Postal 9638, 01065-970 Sao Paulo, SP, Brazil 9 South African Astronomical Observatory, P.O. Box 9, Observatory 7935, Cape Town, South Africa 10 Observatorium Hoher List der Universitätssternwarte Bonn, D-54550 Daun, Germany

The 1994 LBV-like outburst of one of the luminous, hot components of the binary HD 5980 made it the brightest star in the SMC for an interval of 5 months. The most intriguing question to arise in this event is: Why did HD 5980 change from an H-poor WN3 spectrum with veiled OB absorption lines about 20 years ago, to an H-rich WN11 spectrum without central absorption lines during the outburst? In an attempt to answer this enigma, we present and analyze new phasedependent spectroscopic, polarimetric and light-curve observations. Together with other published data, these new observations allow us to considerably improve the orbital parameters, except for the radial velocity amplitudes and hence the masses, which are only roughly constrained.

Especially important in HD 5980 is the strong collision of the two nearly equal pre-outburst winds. The emission-line spectrum generated by the collision tends to mask the underlying line spectra of both components when the system is relatively quiescent. We argue that the pre-erupting system consists of a very luminous but moderately massive, H-rich O-type supergiant, possibly with emission lines, and a low-mass, H-poor, relatively faint WN companion, whose lines are mostly drowned out by wind collision emission, the spectrum of which largely imitates that of a WNE star. It was the O supergiant that erupted in a normal way as an H-rich, visually bright WN11 star. In this way, the need for peculiar evolutionary scenarios (e.g. rapidly from a faint, low-mass, H-poor WNE star to a luminous, H-rich WNL star) is avoided.

Submitted to: The Astrophysical Journal

For preprints, contact moffat@astro.umontreal.ca

Interstellar Abundances in the Magellanic Clouds. I. GHRS Observations of the SMC Star Sk 108

D.E. Welty^{1}, J.T. Lauroesch^{2}, J.C. Blades^{3}, L.M. Hobbs^{4}, and D.G. York^{1}

1 Univ. of Chicago, Dept. of Astronomy & Astrophysics, 5640 S. Ellis Ave., Chicago, IL 60637, USA 2 NASA/Goddard Space Flight Center, Code 631, Greenbelt, MD 20771, USA 3 Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA 4 Univ. of Chicago, Yerkes Observatory, P.O. Box 0258, Williams Bay, WI 53191-0258, USA

We present HST GHRS echelle-B and G160M spectra of the Wolf-Rayet binary Sk 108, located in the northeastern part of the main “bar” of the Small Magellanic Cloud. The spectra show interstellar absorption from C i, O i, Si ii, Si ii*, S ii, P ii, Cr ii, Mn ii, Fe ii, Ni ii, and Zn ii. We find at least 25 interstellar components, which may be grouped via similar kinematics and/or similar relative heavy element abundances into two sets arising in the Galactic disk and halo, plus three sets located in the SMC. The SMC component groups may correspond to large-scale SMC gas complexes identified in H i 21 cm emission surveys.

The relative gas phase abundances for Si ii, Cr ii, Mn ii, Fe ii, Ni ii, and Zn ii found for the SMC interstellar components are similar to those found for interstellar clouds in the Galactic halo. Since the relative total abundances for those elements found for F–G supergiants and gaseous nebulae in the SMC are similar to those present in comparable Galactic objects, we conclude that the interstellar depletion patterns are also similar in the SMC and in the Galactic halo clouds, despite the significantly lower metallicity and dust-to-gas ratio in the SMC. We do not see the distinctive depletion patterns commonly found for cold and warm clouds in the Galactic disk for any of the SMC components along the line of sight to Sk 108.

We discuss some implications of these results for understanding the abundances found for QSO absorption-line systems, and conclude that both nucleosynthetic and depletion effects contribute to the observed abundances – to different degrees for different systems.

Submitted to: The Astrophysical Journal

For preprints, contact welty@oddjob.uchicago.edu

Ultraviolet Color-Magnitude Diagram Studies of Intermediate Age Large Magellanic Cloud Star Clusters I. NGC 1783^{1}

Andrew A. Cole & John S. Gallagher, III^{1} and Wendy L. Freedman & Randy Phelps^{2}

1 Dept. of Astronomy, 5534 Sterling Hall, 475 N. Charter St., University of Wisconsin, Madison, WI 53706, USA 2 Observatories of the Carnegie Institute of Washington, 800 Santa Barbara St., Pasadena, CA 91101, USA

We have imaged the billion-year-old Large Magellanic Cloud star cluster NGC 1783 with the Wide Field/Planetary Camera 2 (WFPC2) in order to obtain broadband photometry at 1600 and 3300 Å. We have identified for the first time the individual stars responsible for the unexpectedly strong ultraviolet light of intermediate-age stellar systems. The behavior of the hot component of aging stellar systems is not well known, and has important implications for the interpretation of light from high-redshift galaxies. Contrary to our expectations, there are far more hot stars present in this cool cluster than can be accounted for by the classical post-asymptotic giant branch. Of particular interest is a group of stars that is strongly concentrated in the core of the cluster. By constructing a relation between the WFPC2 UV colors and stellar effective temperatures, we have found that these stars do not lie in an area of the color-magnitude diagram that is crossed by any theoretical single-star stellar evolution tracks of the cluster’s age. These data are insufficient to unambiguously identify the astrophysical nature of the anomalous UV-bright stars; we discuss possibilities for their origin.

1Based on observations made with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-26555.

Accepted by: The Astronomical Journal For preprints, contact cole@astro.wisc.edu

Obscured AGB stars in the Magellanic Clouds I. IRAS candidates

C. Loup^{1,2}, A.A. Zijlstra^{3}, L.B.F.M. Waters^{4}, M.A.T. Groenewegen^{5}

1 Institut d’Astrophysique de Paris, CNRS, 98 bis Bd. Arago, F–75014 Paris, France 2 European Southern Observatory, Casilla 19001, Santiago 19, Chile 3 European Southern Observatory, Karl–Schwarzschild Strasse 2, D–85748 Garching, Germany 4 Astronomical Institute Anton Pannekoek, University of Amsterdam, Kruislaan 403, NL–1098 SJ Amsterdam, The Netherlands 5 Max–Planck–Institut für Astrophysik, Karl–Schwarzschild Strasse 1, D–85740 Garching, Germany

We have selected 198 IRAS sources in the Large Magellanic Cloud and 11 in the Small Magellanic Cloud that are the best candidates for mass–loosing AGB stars (or possibly post–AGB stars). We used the catalogues of Schwering & Israel (1990) and Reid et al. (1990). They are based on the IRAS pointed observations and have lower detection limits than the Point Source Catalogue. We also made cross–identifications between IRAS sources and optical catalogues.

Our resulting catalogue is divided in 7 tables. Table 1 lists optically known red supergiants and AGB stars for which we found an IRAS counterpart (7 and 52 stars in the SMC and LMC, respectively). Table 2 lists “obscured” (or “cocoon”) AGB stars or late–type supergiants that have been identified as such in previous works through their IRAS counterpart and JHKLM photometry (2 SMC and 34 LMC sources; no optical counterparts). Table 3 lists known planetary nebulae with an IRAS counterpart (4 SMC and 19 LMC PNe). Table 4 lists unidentified IRAS sources that we believe to be good AGB or post–AGB or PNe candidates (11 SMC and 198 LMC sources). Table 5 lists unidentified IRAS sources, which could be any type of object (23 SMC and 121 LMC sources). Table 6 lists IRAS sources associated with foreground stars (29 SMC and 135 LMC stars). Table 7 lists ruled-out IRAS sources associated with H ii regions, hot stars, etc …

We show that the sample of IRAS AGB stars in the Magellanic Clouds is very incomplete. Only AGB stars more luminous than typically 10^{4} L_{⊙} and with a mass-loss rate larger than typically 5·10^{−6} M_{⊙}/yr could be detected by the IRAS satellite. As a consequence, one expects to find very few carbon stars in the IRAS sample. We also expect that most AGB stars with intermediate mass–loss rates have not been discovered yet, neither in optical surveys, nor in the IRAS survey.

Accepted by: Astronomy & Astrophysics Supplement Series For preprints, contact loup@iap.fr

A Catalogue of Compact Radio Sources In and Behind the Large Magellanic Cloud

M. Marx^{1}, J.M. Dickey^{2}, U. Mebold^{1}

1 Radioastronomisches Institut der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2 Department of Astronomy, University of Minnesota, Minneapolis, Minnesota 55455, USA

We present the results of a continuum snapshot survey of a 3^{∘}×4^{∘} region of the Large Magellanic Cloud including the area of the giant molecular cloud and the 30 Doradus nebula. The observations have been carried out with the Australia Telescope Compact Array (ATCA) at 1.4 and 2.4 GHz. Most fields are complete to about 6 mJy peak flux density at 1.4 GHz and to about 3 mJy at 2.4 GHz. The positions, peak and integral flux densities of 113 compact (<54′′) sources detected at 1.4 GHz and of 70 sources (<34′′) detected at 2.4 GHz are presented. Positions are accurate to about 3′′ and peak flux densities are accurate to about 10% or better, depending on the source position relative to the pointing centers. 32 of the sources detected at 1.4 GHz are coincident with Hα objects in the catalogue of Davies, Elliott, and Meaburn; these are possibly intrinsic to the LMC. However, we suppose that most are background objects since the number vs. flux agrees with predictions of extragalactic source counts from other surveys.

Accepted by: Astronomy and Astrophysics

For preprints, contact mmarx@astro.uni-bonn.de

The Superbubble Size Distribution in the Interstellar Medium of Galaxies

M.S. Oey and C.J. Clarke

Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, U.K.

We use the standard, adiabatic shell evolution to predict the differential size distribution N(R) for populations of OB superbubbles in a uniform ISM. Assuming that shell growth stalls upon pressure equilibrium with the ambient ISM, we derive N(R) for simple cases of superbubble creation rate and mechanical luminosity function (MLF). For constant creation and an MLF ϕ(L) ∝ L^{−β}, we find that N(R) ∝ R^{1−2β} for R < R_{e}, and N(R) ∝ R^{4−5β} for R > R_{e}, where the characteristic radius R_{e} ∼ 1300 pc for typical ISM parameters. For R < R_{e}, N(R) is dominated by stalled objects, while for R > R_{e} it is dominated by growing objects. The relation N(R) ∝ R^{1−2β} appears to be quite robust, and also results from momentum-conserving shell evolution. We predict a peak in N(R) corresponding to individual SNRs, and suggest that the contribution of Type Ia SNRs should be apparent in the observed form of N(R). We present expressions for the porosity parameters, Q_{2D} and Q_{3D}, derived from our analysis. Q_{2D} is dominated by the largest superbubbles for β < 2 and individual SNRs for β > 2, whereas Q_{3D} is normally dominated by the few largest shells.

We examine evolutionary effects on the H ii region luminosity function (H ii LF), in order to estimate β. We find that for a nebular luminosity fading with time t, L_{Hα} ∝ t^{−η}, there is a minimum observed slope a_{min} for the Hii LFs. Empirical measurements all show a > a_{min}, therefore implying that usually we may take β = a. We also find that if nebular luminosity is instantaneously extinguished at some given age, rather than continuously fading, no a_{min} will be observed.

Comparison with the largely complete H i hole catalog for the SMC shows surprising agreement in the predicted and observed slope of N(R). This suggests that no other fundamental process is needed to explain the size distribution of shells in the SMC. Further comparison with largely incomplete H i data for M31, M33, and Holmberg II also shows agreement in the slopes, but perhaps hinting at systematic differences between spiral and Im galaxies. We estimate porosities that are substantially < 1 for all of the galaxies except Holmberg II, for which we obtain values ∼ > 1. Most of these galaxies therefore may not be strongly dominated by a hot interstellar component. However, porosity results for the Galaxy remain inconclusive with the available data.

Accepted by: Monthly Notices of the Royal Astronomical Society For preprints, contact oey@ast.cam.ac.uk

Also available from the URL http://www.ast.cam.ac.uk/~oey/oeypubs.html

Abstracts of Non-Refereed Papers

Circumstellar Ring Nebulae Around Early B Supergiants

You-Hua Chu^{1}, Wolfgang Brandner^{1}, and Eva K. Grebel^{2}

1 Astronomy Department, University of Illinois, 1002 W. Green Street, Urbana, IL 61801 2 Astronomisches Institut der Universität Würzburg, Am Hubland, D-97074 Würzburg, Germany

The rings around SN 1987A were shaped by the B3I progenitor Sk−69^{∘}202 and ionized by the UV flash of the supernova. No ring nebulae around B supergiants were known prior to the discovery of SN 1987A’s rings. To determine the frequency of ring nebulae around B supergiants and to search for living counterparts of Sk−69^{∘}202, we have made three pilot searches for ring nebulae around B supergiants in the Galaxy, in 30 Doradus, and in the Small Magellanic Cloud. All three searches yielded negative results, confirming that ring nebulae around B supergiants are rarely seen. We have serendipitously found a circumstellar ring around the B1.5I star Sher 25 in the giant H ii region NGC 3603 in the Galaxy. The similarities between this ring and SN 1987A’s inner ring suggest that Sher 25 is a Galactic counterpart of Sk−69^{∘}202.

To appear in: CTIO/ESO/LCO Workshop “SN 1987A: 10 Years After” For preprints, contact chu@astro.uiuc.edu

Also available from the URL http://www.astro.uiuc.edu/~chu/preprints/BSG.ps

Job Advertisements

Graduate Student Fellowships at the Universities of Bonn and Bochum

The Graduiertenkolleg “The Magellanic System and Other Dwarf Galaxies” of the Universities of Bochum and Bonn invites applications for two graduate student fellowships leading to a PhD in astronomy. Depending on the research pursued, the fellowships will be held either at Bochum or Bonn University.

The successful applicants will find themselves in a pleasant scientific environment, closely collaborating with other graduate students on related topics. They should be interested in galaxy research, in particular in the investigation of dwarf galaxies. Members of the Graduiertenkolleg currently pursue both theoretical and observational studies in the field of galaxy research, and regularly obtain access to major radio and optical telescopes around the world.

The main goal of the Graduiertenkolleg is to furnish comprehensive scientific experience and education of graduate students in astrophysics while working in a team. The scientific focus is on star formation processes and stellar evolution, the interstellar medium, structure and kinematics of the Magellanic Clouds and of more distant dwarf galaxies.

One fellowship is available immediately and the other one on September 1, 1997. The appointment is to the end of 1998 with an expected extension up to a maximum of three years depending on funding and performance. The monthly tax-free salary is DM 1400, plus DM 200 for expenses, plus a DM 300 allowance for married fellows.

Applications along with a curriculum vitae, a transcript of the diploma or master degree certificate, and a short exposé of the diploma or master thesis should be sent to the

Spokesman of the Graduiertenkolleg

Prof. Dr. Uli Klein

Radioastronomisches Institut der Universität Bonn

Auf dem Hügel 71

D-53121 Bonn

Germany

E-Mail: p213ukl@astro.uni-bonn.de

Please note that applicants from abroad will be required to supply a certificate documenting a comprehensive working knowledge of German to be admitted to the PhD program. Alternatively, German classes will have to be taken at either Bochum or Bonn University.

For more information on the Graduiertenkolleg, please consult the URL

                        http://www.astro.uni-bonn.de/~webgk/

Meeting Announcements

Workshop on B[e] stars

June 9–12, 1997 Observatoire de Paris / Institut d’Astrophysique de Paris, France

Since the spectra of B[e] stars are characterized by a large extinction of several magnitudes in the UV and the visual, they have been observed rather infrequently by spectroscopists in the past. As a result, our knowledge of the group is rather incomplete. Although about one hundred galactic objects have been classified as such, only some fifteen have been studied in detail. Besides this, the evolutionary status of the objects is rather controversial; are they pre-main sequence stars, or, on the contrary, are they stars on their way to becoming planetary nebulae?

B[e] stars are characterized by a large infrared excess due to their dust envelope and will thus be easily detectable through new infrared surveys like DENIS, which shall reach K=14.5. Besides the DENIS survey in the south we shall also have the 2MASS data from the northern hemisphere, and more data coming from the micron, the millimeter and the centimeter wavelength regions. Before these new data arrive, it seems that time is ripe to make a survey of our knowledge of these objects.

Although the emphasis of the meeting shall be placed on the spectroscopic side, other aspects will be covered well. Besides the controversial evolutionary status already mentioned, another important point is the relation of B[e] stars with other stellar groups, like Be, Herbig Be/Ae, luminous blue variables and the B[e] stars observed in the Magellanic Clouds.

Preliminary Program:

• Current definition of B[e] stars. Comparison of galactic and extragalactic (SMC and LMC) B[e] stars. Similarities and differences to Herbig Be/Ae, “normal” Be stars, symbiotic objects.

• Distribution of B[e]’s in our Galaxy and a review of global properties like absolute magnitude, distances, extinction, radial velocities and proper motions.

• Spectroscopic observations in various wavelength ranges (X-rays, UV, optical, IR, mm and radio) and variability.

• Photometry, polarimetry and interferometry. Variability.

• Modeling of B[e]’s with dusty environments (winds, discs, bipolar flows).

• B[e]’s and related objects in the evolutionary sequence:

• pre-main-sequence, post-main-sequence?

• Conclusions and redefinition of the group.

Scientific Organizing Committee:

J. Cassinelli (USA), L. Houziaux (Belgium), A.M. Hubert (France), C. Jaschek (France), H. Lamers (The Netherlands), J.M. Marlborough (Canada), A. Miroshnichenko (Russia).

Invited Speakers:

J. Bjorkman, J. Cassinelli, C. Jaschek, H. Lamers, N. Langer, A. Miroshnichenko, S. Owocki, L. Waters, F.Z. Zickgraf, J. Zorec.

Contact: Anne-Marie Hubert, Observatoire de Paris

                            e-mail: dustyBe.coll@obspm.fr